Can you answer this: The length of a side of a rectangle is 4 inches more than the width. If the perimeter is 62 inches, what are the dimensions.
-Thanks

Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

2L+2W=PERIMETER
L=W+4
2(W+4)+2W=62
2W+8+2W=62
4W=62-8
4W=54
W=54/4
W=13.5 in. IS THE WIDTH.
L=13.5+4=17.5 in. IS THE LENGTH.
PROOF:
2*17.5+2*13.5=62
35+27=62
62=62

Answer by rfer(16322)

p=2L+2W
2(x+4)+2x=62
2x+8+2x=62
4x=54
x=13.5
---------------
2(13.5)+2(13.5+4)=62
27+35=62